# Development first Performance evaluation of multi element monolithic HPGe detector for spectroscopy applications

### v1

@@ 5 Conclusion Discussion
We have shown that the detector performs reliably across a wide energy range (20–50 keV) and remains stable under high input count rates. Surface scans with a micro-focused beam confirmed excellent spatial uniformity in most regions, while a few defective areas, which we identified as shallow bulk defects later during the beam tests, showed reduced or no charge collection. A simple suppression model allowed us to estimate the depth and severity of these defects, confirming they are located close to the sensor surface and are likely linked to fabrication/handling issues. Future sensor production will include better quality control (ex, leakage current and capacitance checks) to identify such defects on an earlier stage.
The detector’s performance improved significantly with the optimized electronics housing, particularly in reducing noise and enhancing energy resolution. For example, during recent tests at 5.9 keV (Mn-K), we observed a 30–40 eV improvement in energy resolution after modifications to the electronics box. Currently, the bottleneck is at the sensor level, as the best resolution we can achieve with the best electronic configuration is 248 eV.

### v2

@@ 5 Conclusion Discussion
We have shown that the detector performs reliably across a wide energy range (6–50 keV) and remains stable under high input count rates. Future beam tests will explore performance at higher energies up to 100 keV. Surface scans with a micro-focused beam confirmed excellent spatial uniformity in most regions, while a few defective areas, which we identified as shallow bulk defects later during the beam tests, showed reduced or no charge collection. A simple suppression model allowed us to estimate the depth and severity of these defects, confirming they are located close to the sensor surface and are likely linked to fabrication/handling issues. Future sensor production will include better quality control (ex, leakage current and capacitance checks) to identify such defects on an earlier stage.
The detector’s performance improved significantly with the optimized electronics housing, particularly in reducing noise and enhancing energy resolution. For example, during recent tests at 5.89 keV (Mn-Kα1), we observed a 30–40 eV improvement in energy resolution after modifications to the electronics box. Currently, the bottleneck is at the sensor level, as the best resolution we can achieve with an improved electronic configuration is 248 eV FWHM.
@@ Acknowledgement
A very special thanks to Michele Bordessoule and Martin Chauvin for their valuable guidance and fruitful discussions.